OpenStudy (anonymous):

A man pays $8 at the box office for 3 adults admission tickets and 4 children admission tickets. If children pay half the admission fee charged for the adults, then fee for the adults must be. (A) .80 (B) $1 (C) $1.60 (D) $1.80 (E) $3.20 Is the equation: 3x+4x/2=8?

OpenStudy (anonymous):

Yeah, you got it right, good job. You let the quantity you're looking for, the price of the adult tickets, be x. Then, the cost of 3 of those is 3x, and the cost of 4 child tickets is 4x/2, just like you said, because child tickets are half the cost.

OpenStudy (anonymous):

So, do you see how to solve that equation and get the right answer?

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

So, first simplify 4x/2: that's just 2x. Then rewrite your equation: 3x + 2x = 8 Combine like terms 5x = 8 And isolate x, by dividing by 5 on both sides x = 8/5 = 1.6 And that's C.

OpenStudy (anonymous):

Oh...

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Yerp. But is 1.6 the adult fair, right?

OpenStudy (anonymous):

Right, because that's what x was. You could find a different equation where x is the child fare, and then you'd solve for that instead.

OpenStudy (anonymous):

Oh, okay. Just checking ^^ Thank you.

OpenStudy (anonymous):

No problem, I'm around if you need anything else.

OpenStudy (anonymous):

Oki!